# Provably Safe Reinforcement Learning using Entropy Regularizer

### v2

@@ 6 Conclusion and Future Works
Note that if $w=\frac{T_{max}}{p-p^{s}}$, then $\check{c}_{k}(x,a)=\tilde{c}_{k}(x,a)$, $\forall(x,a)\in\mathcal{X}\times\mathcal{A}$. Hence, optimization (20) becomes equivalent to (11)-(12) and ${V}^{\check{P}_{k}}_{\check{\pi}_{k}}(\check{c}_{k})={V}^{\tilde{P}_{k}}_{\pi_{k}}(\tilde{c}_{k})$, where $(\check{\pi}_{k},\check{P}_{k})$ is the optimal solution of (20) and $(\pi_{k},\tilde{P}_{k})$ is the optimal solution of (7)-(8). In the sequel, we will show that if $w\geq\frac{T_{max}}{p-p_{s}}$, ${V}^{\check{P}_{k}}_{\check{\pi}_{k}}(\check{c}_{k},x_{0})\leq V^{{P}}_{\pi^{*}}(c,x_{0})$.
With the true transition probability kernel $P$, suppose $(\gamma^{s}(x,a)$, $\beta^{s}(x,a)$) and $(\gamma^{*}(x,a)$, $\beta^{*}(x,a)$) be the occupation measure with respect to the safe baseline policy $\pi^{s}$ and the optimal policy $\pi^{*}$, respectively. Consider another occupation measure $(\gamma^{m}(x,a)$, $\beta^{m}(x,a)$) given by: $\gamma^{m}(x,a)=(1-\lambda)\gamma^{s}(x,a)+\lambda\gamma^{*}(x,a)$ and $\beta^{m}(x,a)=(1-\lambda)\beta^{s}(x,a)+\lambda\beta^{*}(x,a)$, where $l_{k}\in[0,1]$. Suppose, $\pi^{m}$ is the policy corresponding to the occupation measures $(\gamma^{m}(x,a)$, $\beta^{m}(x,a)$). We derive a condition such that $(\pi^{m},P)$ is a feasible solution of (20). Since the safety function ${S}^{{P}}_{\pi}(x_{0})$ is a linear function of the occupation measure $\gamma(x,a)$, hence
$\begin{split}&\tilde{S}^{{P}}_{\pi^{m}}(x_{0})\\ &=(1-\lambda)\tilde{S}^{{P}}_{\pi^{s}}(x_{0})+\lambda\tilde{S}^{{P}}_{\pi^{*}}(x_{0})\\ &\overset{(a)}{{=}}(1-\lambda)\mathds{E}_{\pi^{s}}^{x_{0}}\left[\sum_{t=0}^{\tau-1}\big(\tilde{\kappa}(X_{t},A_{t})+3\epsilon_{k}(X_{t},A_{t})\big);P\right]\\ &+\lambda\mathds{E}_{\pi^{*}}^{x_{0}}\left[\sum_{t=0}^{\tau-1}\big(\tilde{\kappa}(X_{t},A_{t})+3\epsilon_{k}(X_{t},A_{t})\big);P\right]\\ &\overset{(b)}{{=}}(1-\lambda)\mathds{E}_{\pi^{s}}^{x_{0}}\left[\sum_{t=0}^{\tau-1}\big({\kappa}(X_{t},A_{t})+4\epsilon_{k}(X_{t},A_{t})\big);P\right]\\ &+\lambda\mathds{E}_{\pi^{*}}^{x_{0}}\left[\sum_{t=0}^{\tau-1}\big({\kappa}(X_{t},A_{t})+4\epsilon_{k}(X_{t},A_{t})\big);P\right]\\ &\overset{(c)}{{\leq}}(1-\lambda)\big({S}^{{P}}_{\pi^{s}}(x_{0})+4\theta_{k}^{\pi_{s}}\big)+\lambda\big({S}^{{P}}_{\pi^{*}}(x_{0})+4\theta_{k}^{\pi^{*}}\big)\\ &\overset{(d)}{{\leq}}(1-\lambda)\big({p}_{s}+4\theta_{k}^{\pi_{s}}\big)+\lambda\big({p}+4\theta_{k}^{\pi^{*}}\big)\end{split}$ (21)
where, $(a)$ is from the definition of $\tilde{S}^{{P}}_{\pi}(x_{0})$ from (20), $(b)$ follows from Lemma (15), $(c)$ follows from the definition of ${S}^{{P}}_{\pi}(x_{0})$ and $\theta_{k}^{\pi}:=\mathds{E}_{\pi}^{x_{0}}\left[\sum_{t=0}^{\tau_{E}-1}\hat{\epsilon}_{k}(X_{t},A_{t});{P}\right]$, $(d)$ is due to the fact that ${S}^{{P}}_{\pi^{s}}(x_{0})\leq{p}_{s}$ and ${S}^{{P}}_{\pi^{*}}(x_{0})\leq{p}$.
@@
$\begin{split}&(1-\lambda)\big({p}_{s}+4\theta_{k}^{\pi_{s}}\big)+\lambda\big({p}+4\theta_{k}^{\pi^{*}}\big)\leq{p}\\ &\implies\lambda\leq\frac{p-p^{s}-4\theta_{k}^{\pi^{s}}}{p-p^{s}-4\theta_{k}^{\pi^{s}}+4\theta_{k}^{\pi^{*}}}\end{split}$ (22)
@@
$\begin{split}&{V}^{P}_{{\pi}^{m}}(\check{c}_{k},x_{0})\leq{V}^{P}_{{\pi}^{*}}({c},x_{0})\\ &\overset{a}{\implies}(1-\lambda){V}^{P}_{{\pi}^{s}}(\check{c}_{k},x_{0})+\lambda{V}^{P}_{{\pi}^{*}}(\check{c}_{k},x_{0})\leq{V}^{P}_{{\pi}^{*}}({c},x_{0})\\ &\overset{b}{\implies}(1-\lambda)\big({V}^{P}_{{\pi}^{s}}({c},x_{0})-4w\theta_{k}^{\pi^{s}}\big)+\lambda\big({V}^{P}_{{\pi}^{*}}({c},x_{0})-4w\theta_{k}^{\pi^{*}}\big)\\ &\hskip 170.71652pt\leq{V}^{P}_{{\pi}^{*}}({c},x_{0})\\ &{\implies}w\big(4\lambda\theta_{k}^{\pi^{*}}+4(1-\lambda)\theta_{k}^{\pi^{s}}\big)\geq(1-\lambda)\big({V}^{P}_{{\pi}^{s}}({c},x_{0})-{V}^{P}_{{\pi}^{*}}({c},x_{0})\big)\\ &{\implies}w\geq\frac{{V}^{P}_{{\pi}^{s}}({c},x_{0})-{V}^{P}_{{\pi}^{*}}({c},x_{0})}{4\theta_{k}^{\pi^{*}}+4\frac{\lambda}{(1-\lambda)}\theta_{k}^{\pi^{s}}}\end{split}$
Relation $(a)$ follows as ${V}^{P}_{{\pi}}({c},x_{0})$ is a linear function of the occupation measure, $(b)$ follows from the definition of $\check{c}_{k}$ and $\theta_{k}^{\pi}$. Since ${V}^{P}_{{\pi}^{s}}({c},x_{0})-{V}^{P}_{{\pi}^{*}}({c},x_{0})\leq T_{max}$ and $4\theta_{k}^{\pi^{*}}+4\frac{\lambda}{(1-\lambda)}\theta_{k}^{\pi^{s}}=p-p^{s}$, we get that $w\geq\frac{T_{max}}{p-p^{s}}\geq\frac{{V}_{{\pi}^{s}}({c},x_{0})-{V}_{{\pi}^{*}}({c},x_{0})}{4\theta_{k}^{\pi^{*}}+4\frac{\lambda}{(1-\lambda)}\theta_{k}^{\pi^{s}}}$. So, if we choose $w=\frac{T_{max}}{p-p_{s}}$ then ${V}^{\check{P}_{k}}_{\check{\pi}_{k}}(\check{c}_{k},x_{0})\leq{V}^{P}_{{\pi}^{m}}(\check{c}_{k},x_{0})\leq{V}^{P}_{{\pi}^{*}}({c},x_{0})$. Hence, ${V}^{\tilde{P}_{k}}_{{\pi}_{k}}(\tilde{c}_{k},x_{0})\leq{V}^{P}_{{\pi}^{*}}({c},x_{0})$.
From the definition of the $c(X^{\tau}_{t},A^{\tau}_{t})$ and $\tilde{c}_{k}(X^{\tau}_{t},A^{\tau}_{t})$, we have that ${V}^{\tilde{P}^{\tau}_{k}}_{{\pi}_{k}}(\tilde{c}_{k},x_{0})={V}^{\tilde{P}_{k}}_{{\pi}_{k}}(\tilde{c}_{k},x_{0})$ and ${V}^{{P^{\tau}}}_{{\pi}^{*}}({c},x_{0})={V}^{{P}}_{{\pi}^{*}}({c},x_{0})$. Thus, we finally have ${V}^{\tilde{P}^{\tau}_{k}}_{{\pi}_{k}}(\tilde{c}_{k},x_{0})\leq{V}^{P^{\tau}_{k}}_{{\pi}^{*}}({c},x_{0})$.
@@
$\begin{split}\sum_{k=1}^{K}\sum_{t=0}^{T_{max}-1}\mathds{E}\big[\frac{1}{\sqrt{{N}^{\tau}_{k}({X}^{\tau}_{k,t},{A}^{\tau}_{k,t})}}\big|{\mathcal{F}}^{\tau}_{k-1}\big]\\ \overset{\overset{}{<}}{\sim}\tilde{\mathcal{O}}\big(\sqrt{|\mathcal{X}||\mathcal{A}|T_{max}K}+|\mathcal{X}||\mathcal{A}|T_{max}\big)\end{split}$
@@
$\sum_{k=1}^{K}\sum_{t=0}^{T_{max}-1}\mathds{E}\big[\frac{1}{{N}^{\tau}_{k}({X}^{\tau}_{k,t},{A}^{\tau}_{k,t})}\big|{\mathcal{F}}^{\tau}_{k-1}\big]\overset{\overset{}{<}}{\sim}\tilde{\mathcal{O}}\big(|\mathcal{X}||\mathcal{A}|T_{max}\big)$
@@
$\begin{split}&\sum_{k=1}^{K}\ \sum_{t,x,a}\sigma_{k,t}(x,a)\sum_{y\in\mathcal{X}}\epsilon^{\tau}_{k}(x,a,y)\\ &\overset{\overset{}{<}}{\sim}\tilde{\mathcal{O}}\left(\sqrt{|\mathcal{X}|^{2}|\mathcal{A}|T_{max}K}+|\mathcal{X}|^{2}|\mathcal{A}|T_{max}\right)\end{split}$
@@
$\begin{split}&\sum_{k=1}^{K}\ \sum_{t,x,a}\sigma_{k,t}(x,a)\sum_{y\in\mathcal{X}}\epsilon^{\tau}_{k}(x,a,y)\\ &=\sum_{k=1}^{K}\ \mathbb{E}^{x_{0}}_{\pi_{k}}\ \left[\sum_{t=0}^{T_{max}-1}\sum_{y\in\mathcal{X}}\epsilon^{\tau}_{k}(X^{\tau}_{k,t},A^{\tau}_{k,t},y);P^{\tau}\right]\\ &\overset{(a)}{\leq}\sum_{k=1}^{K}\mathbb{E}^{x_{0}}_{\pi_{k}}\Big[\sum_{t=0}^{T_{max}-1}\Big(\sqrt{\frac{4L}{N^{\tau}_{k}(X^{\tau}_{k,t},A^{\tau}_{k,t})\vee 1}}\sum_{y\in\mathcal{X}}\sqrt{\hat{P}^{\tau}_{k}(X^{\tau}_{k,t},A^{\tau}_{k,t},y)}\\ &+\sum_{y\in\mathcal{X}}\frac{(14/3)L}{N^{\tau}_{k}(X^{\tau}_{k,t},A^{\tau}_{k,t})\vee 1};P^{\tau}\Big)\Big]\\ &\overset{(b)}{\leq}\sum_{k=1}^{K}\mathbb{E}^{x_{0}}_{\pi_{k}}\Big[2\sqrt{L}\sum_{t=0}^{T_{max}-1}\sqrt{\frac{1}{N^{\tau}_{k}(X^{\tau}_{k,t},A^{\tau}_{k,t})\vee 1}}\\ &\cdot\sqrt{\sum_{y\in\mathcal{X}}\hat{P}^{\tau}_{k}(X^{\tau}_{k,t},A^{\tau}_{k,t},y)}\sqrt{|\mathcal{X}|}\\ &+(14/3)L\sum_{t=0}^{T_{max}-1}\frac{1}{N^{\tau}_{k}(X^{\tau}_{k,t},A^{\tau}_{k,t})\vee 1}|\mathcal{X}|;P^{\tau}\Big]\\ &\overset{\overset{}{<}}{\sim}\tilde{\mathcal{O}}\left(\sqrt{|\mathcal{X}|^{2}|\mathcal{A}||T_{max}|K}+|\mathcal{X}|^{2}|\mathcal{A}|T_{max}\right)\end{split}$
@@
due to Cauchy-Schwarz inequality. The final inequality follows from Lemma 19 and Lemma 20.
@@
$\begin{split}&V^{P^{\tau}}_{\pi}(c,x)-V^{\bar{P}^{\tau}}_{\pi}(\bar{c},x)\\ &=\mathds{E}^{x}_{\pi}\sum_{t=0}^{T_{max}-1}\Big[\left(c(X^{\tau}_{t},A^{\tau}_{t})-\bar{c}(X^{\tau}_{t},A^{\tau}_{t})\right)\\ &+\sum_{y\in\mathcal{X}}\big(P^{\tau}(X^{\tau}_{t},A^{\tau}_{t},y)-\bar{P}^{\tau}(X^{\tau}_{t},A^{\tau}_{t},y)\big)V^{{P}^{\tau}}_{\pi}({c},y);\bar{P}^{\tau}\Big]\\ &=\mathds{E}^{x}_{\pi}\sum_{t=0}^{T_{max}-1}\Big[\left(c(X^{\tau}_{t},A^{\tau}_{t})-\bar{c}(X^{\tau}_{t},A^{\tau}_{t})\right)\\ &+\sum_{y\in\mathcal{X}}\big(P^{\tau}(X^{\tau}_{t},A^{\tau}_{t},y)-\bar{P}^{\tau}(X^{\tau}_{t},A^{\tau}_{t},y)\big)V^{\bar{P}^{\tau}}_{\pi}(\bar{c},y);{P}^{\tau}\Big]\end{split}$
@@
For any $x\in\mathcal{X}$, the following is true for the stopped MDP $M^{\tau}$:
$\begin{split}&V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},x)\leq V^{P^{\tau}}_{\pi_{k}}(c,x)\end{split}$
@@
$\begin{split}&V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},x)-V^{P^{\tau}}_{\pi_{k}}(c,x)\\ &\overset{(a)}{=}\mathds{E}^{x}_{\pi_{k}}\sum_{t=0}^{T_{max}}\big[\left(\tilde{c}_{k}(X^{\tau}_{t},A^{\tau}_{t})-{c}(X^{\tau}_{t},A^{\tau}_{t})\right)\\ &+\sum_{y\in\mathcal{X}}\big(\tilde{P}^{\tau}_{k}(X^{\tau}_{t},A^{\tau}_{t},y)-{P}^{\tau}(X^{\tau}_{t},A^{\tau}_{t},y)\big)V^{{P}^{\tau}}_{\pi}({c},y);\tilde{P}^{\tau}_{k}\big]\\ &\overset{(b)}{\leq}\mathds{E}^{x}_{\pi_{k}}\sum_{t=0}^{T_{max}}\Big[\left({c}(X^{\tau}_{t},A^{\tau}_{t})-\frac{4T_{max}}{p-p^{s}}\hat{\epsilon}^{\tau}_{k}(X^{\tau}_{t},A^{\tau}_{t})-{c}(X^{\tau}_{t},A^{\tau}_{t})\right)\\ &+\sum_{y\in\mathcal{X}}\epsilon^{\tau}_{k}(X^{\tau}_{t},A^{\tau}_{t},y)T_{max};\tilde{P}^{\tau}_{k}\Big]\\ &\overset{}{\leq}\mathds{E}^{x}_{\pi_{k}}\sum_{t=0}^{T_{max}}\Big[-\frac{4T_{max}}{p-p^{s}}\hat{\epsilon}^{\tau}_{k}(X^{\tau}_{t},A^{\tau}_{t})+T_{max}\hat{\epsilon}^{\tau}_{k}(X^{\tau}_{t},A^{\tau}_{t});\tilde{P}^{\tau}_{k}\Big]\\ &\overset{(c)}{\leq}0\end{split}$
Relation $(a)$ follows from Lemma 23. Inequality $(b)$ follows from the definition of $\tilde{c}_{k}(x,a)$, Lemma 15 and due to the fact that $V^{P^{\tau}}_{\pi}(c,x)\leq T_{max}$ for all $x\in\mathcal{X}$. Finally, we get inequality $(c)$ as $\frac{T_{max}}{p-p^{s}}\geq T_{max}$. The following result is in line with Lemma $33$ in [10].

### v3

@@ 6 Conclusion and Future Works
Note that if $w=\frac{T_{max}}{p-p^{s}}$, then $\check{c}_{k}(x,a)=\tilde{c}_{k}(x,a)$, $\forall(x,a)\in\mathcal{X}\times\mathcal{A}$. Hence, optimization (21) becomes equivalent to (7)-(8) and ${V}^{\check{P}_{k}}_{\check{\pi}_{k}}(\check{c}_{k},x_{0})={V}^{\tilde{P}_{k}}_{\pi_{k}}(\tilde{c}_{k},x_{0})$, where $(\check{\pi}_{k},\check{P}_{k})$ is the optimal solution of (21) and $(\pi_{k},\tilde{P}_{k})$ is the optimal solution of (7)-(8). In the sequel, we will show that if $w\geq\frac{T_{max}}{p-p_{s}}$, ${V}^{\check{P}_{k}}_{\check{\pi}_{k}}(\check{c}_{k},x_{0})\leq V^{{P}}_{\pi^{*}}(c,x_{0})$.
With the true transition probability kernel $P$, suppose $(\gamma^{s}(x,a)$, $\beta^{s}(x,a)$) and $(\gamma^{*}(x,a)$, $\beta^{*}(x,a)$) be the occupation measure with respect to the safe baseline policy $\pi^{s}$ and the optimal policy $\pi^{*}$, respectively. Consider another occupation measure $(\gamma^{m}(x,a)$, $\beta^{m}(x,a)$) given by: $\gamma^{m}(x,a)=(1-\lambda)\gamma^{s}(x,a)+\lambda\gamma^{*}(x,a)$ and $\beta^{m}(x,a)=(1-\lambda)\beta^{s}(x,a)+\lambda\beta^{*}(x,a)$, where $\lambda\in[0,1]$. Suppose, $\pi^{m}$ is the policy corresponding to the occupation measures $(\gamma^{m}(x,a)$, $\beta^{m}(x,a)$). We derive a condition such that $(\pi^{m},P)$ is a feasible solution of (21). Since the safety function ${S}^{{P}}_{\pi}(x_{0})$ is a linear function of the occupation measure $\gamma(x,a)$, hence
$\begin{split}&\tilde{S}^{{P}}_{\pi^{m}}(x_{0})\\ &=(1-\lambda)\tilde{S}^{{P}}_{\pi^{s}}(x_{0})+\lambda\tilde{S}^{{P}}_{\pi^{*}}(x_{0})\\ &\overset{(a)}{{=}}(1-\lambda)\mathds{E}_{\pi^{s}}^{x_{0}}\left[\sum_{t=0}^{\tau-1}\big(\hat{\kappa}(X_{t},A_{t})+3\epsilon_{k}(X_{t},A_{t})\big);P\right]\\ &+\lambda\mathds{E}_{\pi^{*}}^{x_{0}}\left[\sum_{t=0}^{\tau-1}\big(\hat{\kappa}(X_{t},A_{t})+3\epsilon_{k}(X_{t},A_{t})\big);P\right]\\ &\overset{(b)}{{\leq}}(1-\lambda)\mathds{E}_{\pi^{s}}^{x_{0}}\left[\sum_{t=0}^{\tau-1}\big({\kappa}(X_{t},A_{t})+4\epsilon_{k}(X_{t},A_{t})\big);P\right]\\ &+\lambda\mathds{E}_{\pi^{*}}^{x_{0}}\left[\sum_{t=0}^{\tau-1}\big({\kappa}(X_{t},A_{t})+4\epsilon_{k}(X_{t},A_{t})\big);P\right]\\ &\overset{(c)}{{\leq}}(1-\lambda)\big({S}^{{P}}_{\pi^{s}}(x_{0})+4\theta_{k}^{\pi_{s}}\big)+\lambda\big({S}^{{P}}_{\pi^{*}}(x_{0})+4\theta_{k}^{\pi^{*}}\big)\\ &\overset{(d)}{{\leq}}(1-\lambda)\big({p}_{s}+4\theta_{k}^{\pi_{s}}\big)+\lambda\big({p}+4\theta_{k}^{\pi^{*}}\big).\end{split}$ (22)
Relation $(a)$ is from the definition of $\tilde{S}^{{P}}_{\pi}(x_{0})$ from (21), $(b)$ follows from Lemma (15), $(c)$ follows from the definition of ${S}^{{P}}_{\pi}(x_{0})$ and $\theta_{k}^{\pi}:=\mathds{E}_{\pi}^{x_{0}}\left[\sum_{t=0}^{\tau_{E}-1}\hat{\epsilon}_{k}(X_{t},A_{t});{P}\right]$, $(d)$ is due to the fact that ${p}_{s}:={S}^{{P}}_{\pi^{s}}(x_{0})$ and ${S}^{{P}}_{\pi^{*}}(x_{0})\leq{p}$.
@@
$\begin{split}&(1-\lambda)\big({p}_{s}+4\theta_{k}^{\pi_{s}}\big)+\lambda\big({p}+4\theta_{k}^{\pi^{*}}\big)\leq{p}\\ &\implies\lambda\leq\frac{p-p^{s}-4\theta_{k}^{\pi^{s}}}{p-p^{s}-4\theta_{k}^{\pi^{s}}+4\theta_{k}^{\pi^{*}}}.\end{split}$ (23)
@@
$\begin{split}&{V}^{P}_{{\pi}^{m}}(\check{c}_{k},x_{0})\leq{V}^{P}_{{\pi}^{*}}({c},x_{0})\\ &\overset{a}{\implies}(1-\lambda){V}^{P}_{{\pi}^{s}}(\check{c}_{k},x_{0})+\lambda{V}^{P}_{{\pi}^{*}}(\check{c}_{k},x_{0})\leq{V}^{P}_{{\pi}^{*}}({c},x_{0})\\ &\overset{b}{\implies}(1-\lambda)\big({V}^{P}_{{\pi}^{s}}({c},x_{0})-4w\theta_{k}^{\pi^{s}}\big)+\lambda\big({V}^{P}_{{\pi}^{*}}({c},x_{0})-4w\theta_{k}^{\pi^{*}}\big)\\ &\hskip 170.71652pt\leq{V}^{P}_{{\pi}^{*}}({c},x_{0})\\ &{\implies}w\big(4\lambda\theta_{k}^{\pi^{*}}+4(1-\lambda)\theta_{k}^{\pi^{s}}\big)\geq(1-\lambda)\big({V}^{P}_{{\pi}^{s}}({c},x_{0})-{V}^{P}_{{\pi}^{*}}({c},x_{0})\big)\\ &{\implies}w\geq\frac{{V}^{P}_{{\pi}^{s}}({c},x_{0})-{V}^{P}_{{\pi}^{*}}({c},x_{0})}{4\theta_{k}^{\pi^{s}}+4\frac{\lambda}{(1-\lambda)}\theta_{k}^{\pi^{*}}}.\end{split}$
The relation $(a)$ follows as ${V}^{P}_{{\pi}}({c},x_{0})$ is a linear function of the occupation measure, $(b)$ follows from the definition of $\check{c}_{k}$ and $\theta_{k}^{\pi}$. If we choose $\lambda=\frac{p-p^{s}-4\theta_{k}^{\pi^{s}}}{p-p^{s}-4\theta_{k}^{\pi^{s}}+4\theta_{k}^{\pi^{*}}}$, then $4\theta_{k}^{\pi^{s}}+4\frac{\lambda}{(1-\lambda)}\theta_{k}^{\pi^{*}}=p-p^{s}$. Now, since ${V}^{P}_{{\pi}^{s}}({c},x_{0})-{V}^{P}_{{\pi}^{*}}({c},x_{0})\leq T_{max}$, we get that $w\geq\frac{T_{max}}{p-p^{s}}\geq\frac{{V}_{{\pi}^{s}}({c},x_{0})-{V}_{{\pi}^{*}}({c},x_{0})}{4\theta_{k}^{\pi^{*}}+4\frac{\lambda}{(1-\lambda)}\theta_{k}^{\pi^{s}}}$. So, if we choose $w=\frac{T_{max}}{p-p_{s}}$ then ${V}^{\check{P}_{k}}_{\check{\pi}_{k}}(\check{c}_{k},x_{0})\leq{V}^{P}_{{\pi}^{m}}(\check{c}_{k},x_{0})\leq{V}^{P}_{{\pi}^{*}}({c},x_{0})$. Hence, ${V}^{\tilde{P}_{k}}_{{\pi}_{k}}(\tilde{c}_{k},x_{0})\leq{V}^{P}_{{\pi}^{*}}({c},x_{0})$.
From the definition of the $c(X^{\tau}_{t},A^{\tau}_{t})$ and $\tilde{c}_{k}(X^{\tau}_{t},A^{\tau}_{t})$, we have that ${V}^{\tilde{P}^{\tau}_{k}}_{{\pi}_{k}}(\tilde{c}_{k},x_{0})={V}^{\tilde{P}_{k}}_{{\pi}_{k}}(\tilde{c}_{k},x_{0})$ and ${V}^{{P^{\tau}}}_{{\pi}^{*}}({c},x_{0})={V}^{{P}}_{{\pi}^{*}}({c},x_{0})$. Thus, we finally have
${V}^{\tilde{P}^{\tau}_{k}}_{{\pi}_{k}}(\tilde{c}_{k},x_{0})\leq{V}^{P^{\tau}_{k}}_{{\pi}^{*}}({c},x_{0}).\qed$
@@
$\begin{split}\sum_{k=1}^{K}\sum_{t=0}^{T_{max}-1}\mathds{E}\big[\frac{1}{\sqrt{{N}^{\tau}_{k}({X}^{\tau}_{k,t},{A}^{\tau}_{k,t})}}\big|{\mathcal{F}}^{\tau}_{k-1}\big]\\ \overset{\overset{}{<}}{\sim}\tilde{\mathcal{O}}\big(\sqrt{|\mathcal{X}||\mathcal{A}|T_{max}K}+|\mathcal{X}||\mathcal{A}|T_{max}\big).\end{split}$
@@
$\sum_{k=1}^{K}\sum_{t=0}^{T_{max}-1}\mathds{E}\big[\frac{1}{{N}^{\tau}_{k}({X}^{\tau}_{k,t},{A}^{\tau}_{k,t})}\big|{\mathcal{F}}^{\tau}_{k-1}\big]\overset{\overset{}{<}}{\sim}\tilde{\mathcal{O}}\big(|\mathcal{X}||\mathcal{A}|T_{max}\big).$
@@
$\begin{split}&\sum_{k=1}^{K}\ \sum_{t,x,a}\sigma_{k,t}(x,a)\sum_{y\in\mathcal{X}}\epsilon^{\tau}_{k}(x,a,y)\\ &\overset{\overset{}{<}}{\sim}\tilde{\mathcal{O}}\left(\sqrt{|\mathcal{X}|^{2}|\mathcal{A}|T_{max}K}+|\mathcal{X}|^{2}|\mathcal{A}|T_{max}\right).\end{split}$
@@
$\begin{split}&\sum_{k=1}^{K}\ \sum_{t,x,a}\sigma_{k,t}(x,a)\sum_{y\in\mathcal{X}}\epsilon^{\tau}_{k}(x,a,y)\\ &=\sum_{k=1}^{K}\ \mathbb{E}^{x_{0}}_{\pi_{k}}\ \left[\sum_{t=0}^{T_{max}-1}\sum_{y\in\mathcal{X}}\epsilon^{\tau}_{k}(X^{\tau}_{k,t},A^{\tau}_{k,t},y);P^{\tau}\right]\\ &\overset{(a)}{\leq}\sum_{k=1}^{K}\mathbb{E}^{x_{0}}_{\pi_{k}}\Big[\sum_{t=0}^{T_{max}-1}\Big(\sqrt{\frac{4L}{N^{\tau}_{k}(X^{\tau}_{k,t},A^{\tau}_{k,t})\vee 1}}\sum_{y\in\mathcal{X}}\sqrt{\hat{P}^{\tau}_{k}(X^{\tau}_{k,t},A^{\tau}_{k,t},y)}\\ &+\sum_{y\in\mathcal{X}}\frac{(14/3)L}{N^{\tau}_{k}(X^{\tau}_{k,t},A^{\tau}_{k,t})\vee 1};P^{\tau}\Big)\Big]\\ &\overset{(b)}{\leq}\sum_{k=1}^{K}\mathbb{E}^{x_{0}}_{\pi_{k}}\Big[2\sqrt{L}\sum_{t=0}^{T_{max}-1}\sqrt{\frac{1}{N^{\tau}_{k}(X^{\tau}_{k,t},A^{\tau}_{k,t})\vee 1}}\\ &\cdot\sqrt{\sum_{y\in\mathcal{X}}\hat{P}^{\tau}_{k}(X^{\tau}_{k,t},A^{\tau}_{k,t},y)}\sqrt{|\mathcal{X}|}\\ &+(14/3)L\sum_{t=0}^{T_{max}-1}\frac{1}{N^{\tau}_{k}(X^{\tau}_{k,t},A^{\tau}_{k,t})\vee 1}|\mathcal{X}|;P^{\tau}\Big]\\ &\overset{\overset{}{<}}{\sim}\tilde{\mathcal{O}}\left(\sqrt{|\mathcal{X}|^{2}|\mathcal{A}||T_{max}|K}+|\mathcal{X}|^{2}|\mathcal{A}|T_{max}\right).\end{split}$
@@
due to the Cauchy-Schwarz inequality. The final inequality follows from Lemma 19 and Lemma 20. ∎
@@
$\begin{split}&V^{P^{\tau}}_{\pi}(c,x)-V^{\bar{P}^{\tau}}_{\pi}(\bar{c},x)\\ &=\mathds{E}^{x}_{\pi}\sum_{t=0}^{T_{max}-1}\Big[\left(c(X^{\tau}_{t},A^{\tau}_{t})-\bar{c}(X^{\tau}_{t},A^{\tau}_{t})\right)\\ &+\sum_{y\in\mathcal{X}}\big(P^{\tau}(X^{\tau}_{t},A^{\tau}_{t},y)-\bar{P}^{\tau}(X^{\tau}_{t},A^{\tau}_{t},y)\big)V^{{P}^{\tau}}_{\pi}({c},y);\bar{P}^{\tau}\Big]\\ &=\mathds{E}^{x}_{\pi}\sum_{t=0}^{T_{max}-1}\Big[\left(c(X^{\tau}_{t},A^{\tau}_{t})-\bar{c}(X^{\tau}_{t},A^{\tau}_{t})\right)\\ &+\sum_{y\in\mathcal{X}}\big(P^{\tau}(X^{\tau}_{t},A^{\tau}_{t},y)-\bar{P}^{\tau}(X^{\tau}_{t},A^{\tau}_{t},y)\big)V^{\bar{P}^{\tau}}_{\pi}(\bar{c},y);{P}^{\tau}\Big].\end{split}$
@@
For any $x\in\mathcal{X}$, the following is true for the stopped MDP $M^{\tau}$ within the event $\mathcal{E}$:
$\begin{split}&V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},x)\leq V^{P^{\tau}}_{\pi_{k}}(c,x).\end{split}$
@@
$\begin{split}&V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},x)-V^{P^{\tau}}_{\pi_{k}}(c,x)\\ &\overset{(a)}{=}\mathds{E}^{x}_{\pi_{k}}\sum_{t=0}^{T_{max}-1}\big[\left(\tilde{c}_{k}(X^{\tau}_{t},A^{\tau}_{t})-{c}(X^{\tau}_{t},A^{\tau}_{t})\right)\\ &+\sum_{y\in\mathcal{X}}\big(\tilde{P}^{\tau}_{k}(X^{\tau}_{t},A^{\tau}_{t},y)-{P}^{\tau}(X^{\tau}_{t},A^{\tau}_{t},y)\big)V^{{P}^{\tau}}_{\pi}({c},y);\tilde{P}^{\tau}_{k}\big]\\ &\overset{(b)}{\leq}\mathds{E}^{x}_{\pi_{k}}\sum_{t=0}^{T_{max}-1}\Big[\left({c}(X^{\tau}_{t},A^{\tau}_{t})-\frac{4T_{max}}{p-p^{s}}\hat{\epsilon}^{\tau}_{k}(X^{\tau}_{t},A^{\tau}_{t})-{c}(X^{\tau}_{t},A^{\tau}_{t})\right)\\ &+\sum_{y\in\mathcal{X}}\epsilon^{\tau}_{k}(X^{\tau}_{t},A^{\tau}_{t},y)T_{max};\tilde{P}^{\tau}_{k}\Big]\\ &\overset{}{\leq}\mathds{E}^{x}_{\pi_{k}}\sum_{t=0}^{T_{max}-1}\Big[-\frac{4T_{max}}{p-p^{s}}\hat{\epsilon}^{\tau}_{k}(X^{\tau}_{t},A^{\tau}_{t})+T_{max}\hat{\epsilon}^{\tau}_{k}(X^{\tau}_{t},A^{\tau}_{t});\tilde{P}^{\tau}_{k}\Big]\\ &\overset{(c)}{\leq}0.\end{split}$
Relation $(a)$ follows from Lemma 23. Inequality $(b)$ follows from the definition of $\tilde{c}_{k}(x,a)$, Lemma 15 and due to the fact that $V^{P^{\tau}}_{\pi}(c,x)\leq T_{max}$ for all $x\in\mathcal{X}$. Finally, we get inequality $(c)$ as $\frac{T_{max}}{p-p^{s}}\geq T_{max}$. ∎ The following result is in line with Lemma $33$ in [10].